OpenStudy (anonymous):

4x=y^2 determine if the equation defines y as a function of x

OpenStudy (anonymous):

i have no clue bout this

OpenStudy (anonymous):

Suppose \(x=1\). Then you have \[4=y^2\] There are two possible solutions here: \(y=2\) and \(y=-2\). Since there are two values of \(y\) for one value of \(x\), \(y\) is not a function of \(x\).

OpenStudy (anonymous):

y did we suppose x is one?

OpenStudy (anonymous):

can we suppose x is any number but keep it as 1 to keep it easy?

OpenStudy (anonymous):

You could really let \(x\) be anything, within reason. In general, \(x\) can't be negative because \(y^2>0\) for any \(y\). So you're left with \(x\ge0\). If you let \(x=0\), then you have \(0=y^2~\Rightarrow~y=0\) (one \(y\) for one \(x\)). But if you let \(x>0\), then you have \(x=y^2~\Rightarrow~y=\pm\sqrt x\) (one \(y\) for more than one \(x\)). And yes, I chose 1 for convenience.

OpenStudy (anonymous):

thanx buddy

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

see ya later!!!

OpenStudy (anonymous):

Bye